### v1

@@ 2.2 Image Registration
Since the body shape of the patients possibly changes in the course of therapy, an affine registration is performed that also allows for shearing and scaling in addition to the rotation and translation of a rigid transformation. The follow-up CT scan is registered towards the baseline CT scan. As a result we obtain a shear and rotation matrix $\mathbf{A}=(a_{11},...\ a_{33})$ and the translation vector $\mathbf{t}=(t_{1},t_{2},t_{3})$, along with the center of rotation $\mathbf{c}=(c_{1},c_{2},c_{3})$, defined as the center of the baseline scan[ElastixManual]. The transformation is then applied on the follow-up PET scan as well as on its 3D bone segmentation mask.
@@ 4 Discussion
Comparing the result of our AI PET/CT lesion tracking with an accuracy of 83.3% to other CNN approaches for lesion tracking in CT scans, we reach a similar performance level, even tough in our cases have the additional challenge of a high number of closely neighboring bone lesions that have to be distinguished. For a lower number of neighboring lesions, Hering et al. [22] achieve 80% test accuracy during whole-body soft-tissue lesion tracking and Rafael-Palou et al. [25] re-identify 88.8% of pulmonary nodules in follow-up scans. Within the group of remaining lesions, we even reach re-identification rate of 89.3%.
@@
The fact that we used a heterogeneous dataset for the training and testing of our pipeline, including images acquired by different scanners and with even different slice thicknesses in z-direction shows the robustness of the approach. It is capable of handling clinical data. As our developed principle of AI lesion tracking is independent of bone lesions in PSMA-PET/CT scans, it can also be applied on other types of PET/CT scans.
Those were only an application example.
Our findings are encouraging for the use in multimodal imaging. Nevertheless, the following limitations should be noted. In our approach, inaccuracies in the segmentation by the qPSMA software[27] are forwarded to the tracking result. Especially, in cases of disease progress and growing lesions, the segmentation software wrongfully draws large, fused follow-up lesions spreading in several bones instead of segmenting several uptake foci separately. In this case, the completely separated lesion segmentation and tracking causes errors. This is still a challenge for our pipeline and lowers the detection rate of fused lesion cases. Due to the segmentation method [27] being specialized to the segmentation of bone lesions in PSMA-PET/CT scans, we chose those as an application example. However, a new and more pipeline-optimized segmentation approach is currently in development. In addition, our test set contained a high number of resolved baseline and new follow-up lesions right beside remaining lesion pairs. Even tough our pipeline handles most of the neighboring lesions well, this leads to a lower success rate for resolved and new lesion detection, which is to be improved in future works. A possibility could be to integrate a transformer network structure into our pipeline, as proposed by Tang et al.[40]. The network might increase the anatomical precision for the described cases. In this context, it has be to noted that for the closely located lesions human readers reach their limits, too, concerning the differentiation of split and new, or resolved and fused lesions. For this reason, the gold standard is far from being perfect. This is demonstrated by figure 9C, where a CNN is outperforming a nuclear physician’s decision, proving the ground truth dataset wrong in this case.
@@ 5 Conclusion
M.S. is currently employed by Siemens Healthineers but was not affiliated with the company during his contributions to the research phase of this manuscript.

### v3

@@ 2.2 Image Registration
Since the body shape of the patients possibly changes in the course of therapy, an affine registration is performed in the pipeline that also allows for shearing and scaling in addition to the rotation and translation of a rigid transformation. The follow-up CT scan is registered towards the baseline CT scan. As a result we obtain a shear and rotation matrix $\mathbf{A}=(a_{11},...\ a_{33})$ and the translation vector $\mathbf{t}=(t_{1},t_{2},t_{3})$, along with the center of rotation $\mathbf{c}=(c_{1},c_{2},c_{3})$, defined as the center of the baseline scan. The transformation is then applied on the follow-up PET scan as well as on its 3D bone segmentation mask.
@@ 4 Discussion
Comparing the result of our AI PET/CT lesion tracking with an accuracy of 83.3% to other CNN approaches for lesion tracking in CT scans, we reach a similar performance level, even though our cases have the additional challenge of a high number of closely neighboring bone lesions that have to be distinguished. For a lower number of neighboring lesions, Hering et al. [22] achieve 80% test accuracy during whole-body soft-tissue lesion tracking and Rafael-Palou et al. [25] re-identify 88.8% of pulmonary nodules in follow-up scans. However, within the group of remaining lesions, we even reach a re-identification rate of 89.3%.
@@
The fact that we used a heterogeneous dataset for the training and testing of our pipeline, including images acquired by different scanners and with even different slice thicknesses in z-direction (sec. 3.1) shows the robustness of the approach. It is capable of handling clinical data. As our developed principle of AI lesion tracking is independent of bone lesions in PSMA-PET/CT scans, it can also be applied on other types of PET/CT scans. In fact, PSMA-PET/CT scans have only been chosen as a difficult application example due to the numerous bone lesions.
Our findings are encouraging for the use in multimodal imaging. Nevertheless, the following limitations should be noted. In our approach, inaccuracies in the segmentation by the qPSMA software[27] are forwarded to the tracking result. Especially, in cases of disease progress and growing lesions, the segmentation software partially wrongfully draws large, fused follow-up lesions spreading in several bones instead of segmenting several uptake foci separately. In this case, the completely separated lesion segmentation and tracking causes errors. This is still a challenge for our pipeline and lowers the detection rate of fused lesion cases. The used segmentation method [27] was chosen due to its specialization on segmentation of bone lesions in PSMA-PET/CT scans. However, a new and more pipeline-optimized segmentation approach is currently in development. In addition, our test set contained a high number of resolved baseline and new follow-up lesions right beside remaining lesion pairs. Even though our pipeline handles most of the neighboring lesions well, this leads to a lower success rate for resolved and new lesion detection, which is to be improved in future works. A possibility could be to integrate a transformer network structure into our pipeline, as proposed by Tang et al.[40]. The network might increase the anatomical precision for the described cases. In this context, it has be to noted that for the closely located lesions human readers reach their limits, too, concerning the differentiation of split and new, or resolved and fused lesions. For this reason, the gold standard is far from being perfect. This is demonstrated by fig. 9C, where a CNN is outperforming a nuclear physician’s decision, proving the ground truth dataset wrong in this case.